# LSAST: Enhancing Cybersecurity through LLM-supported Static Application Security Testing

### v1

@@ Abstract
In the fast-evolving landscape of cyber security, Large Language Models (LLM’s) play a pivotal role, continually improving their ability to analyze software code. This paper introduces a novel approach to vulnerability scanning by integrating conservative SAST (Static Application Security Testing) scanners with LLM capabilities, resulting in the creation of LSAST (LLM-supported Static Application Security Testing). Our approach significantly enhances the performance of LLM’s in vulnerability scanning, establishing a new standard in this field. We benchmark LSAST’s efficiency and compare its results with a state-of-the-art LLM. Additionally, we address the inherent drawbacks of LLM’s in vulnerability scanning: their reliance on static training datasets, which leads to the exclusion of the latest vulnerabilities, and the privacy concerns associated with sending code to third-party LLM providers. To mitigate these issues, we utilize a open-source LLM to ensure privacy and employ a novel approach to gather relevant vulnerability information, thereby equipping the LLM with up-to-date knowledge.
@@ I Introduction
The realm of cyber security is characterized by its dynamic nature, evolving rapidly alongside advancements in Large Language Models (LLM’s). These sophisticated models continuously enhance their ability to analyze and comprehend software code, due to their access to comprehensive training data, which includes both code and vulnerability information.
Using large language models (LLM’s) for vulnerability detection in code has already shown promising results. In this paper, we aim to enhance the capabilities of LLM’s in identifying vulnerabilities in software code. Our approach combines conservative Static Application Security Testing (SAST) scanners, such as Bearer, with LLM’s and thereby achieves synergy in code scanning. We accomplish this by engineering a prompt that includes the results of a conservative SAST scanner and instructing the LLM to search only for new vulnerabilities not identified by the SAST scanner. This method allows the LLM’s task to be strongly supported by the SAST scanner results, leading to higher performance in identifying additional vulnerabilities. Our findings indicate that this approach significantly enhances the LLM’s vulnerability detection capabilities and improves the quality of the results, making this method crucial for future use of LLM’s in SAST scanning.
However, a notable limitation of LLM’s is their reliance on a static training dataset, leading to the exclusion of the latest vulnerabilities. This creates significant security risks, when utilizing LLM’s for Vulnerability Scanning, as vulnerability scanners rely on access to the most current vulnerabilities to adequately address emerging threats.
@@
Furthermore, utilizing LLM’s for vulnerability scanning introduces additional challenges, notably when transmitting code to API endpoints of third-party LLM providers. This practice exposes sensitive code, opening the door to potential privacy and security risks, notably unauthorized access by malicious entities. Additionally, there’s the concern of the code being repurposed for future model training, creating a vulnerability to training data extraction attacks. Addressing these concerns is necessary to prevent malicious exploitation of possible vulnerabilities in the code.
In this paper, we aim to tackle these challenges head-on. Our approach centers on utilizing a locally hostable instance of Llama-3, to mitigate security and privacy risks associated with third-party LLM providers. Remarkably, existing research has acknowledged these privacy and security issues but has yet to benchmark open-source LLM’s for Vulnerability Scans. Therefore, our paper represents the first attempt to showcase the efficiency of a locally hostable LLM-Scanner in identifying vulnerabilities.
To ensure LLM’s have access to the most up-to-date vulnerability knowledge, we implement a state-of-the-art Knowledge Retrieval system. This involves aggregating the latest vulnerability reports from HackerOne, a prominent Bug Bounty platform known for its vast repository of reports submitted by ethical hackers worldwide. Leveraging the Retrieval-Augmented Generation Framework, we enable LLM’s to acquire knowledge of the most recent vulnerabilities.
By addressing privacy and security concerns, integrating current vulnerability knowledge into LLM’s, and enhancing results through SAST scanner outputs, we introduce LSAST (LLM-supported Static Application Security Scanning), heralding a new era in static analysis within IT security.
Our research further aims to benchmark LSAST’s efficiency by comparing its detection capabilities to those of unenhanced LLMs. Additionally, we evaluate LSAST’s performance in real-world scenarios by scanning a Python package used in industry. This analysis demonstrates that LSAST can identify vulnerabilities overlooked by conventional scanners, underscoring its importance as an integral component of comprehensive security scanning setups.
@@ II Background
II-A Overview
With the global cost of cybercrime projected to surge from \$8.15 trillion to \$13.82 trillion [6] and hackers launching an average of 26,000 attacks daily, or one every three seconds [3], it is clear that cybersecurity is more critical than ever in the digital age.
As the IT landscape expands, cyber threats have evolved from simple attacks to sophisticated and persistent threats. To counter these threats, a range of cybersecurity measures must be implemented.
One widely adopted measure is security scanning, which plays a crucial role in identifying potential risks within applications or networks.
II-B Different kinds of security scanners
Predominantly, security scanning can be classified into two categories: DAST (Dynamic Application Security Testing) and SAST (Static Application Security Testing).
@@ II-B Different kinds of security scanners
DAST scanners treat the system or application as a black-box, meaning they operate without knowledge of the internal workings. They identify weaknesses by interacting with the system or application during runtime. By sending specially crafted requests and analyzing the responses, DAST scanners mimic threat actors attempting various types of attacks, such as SQL Injection, Cross-Site Scripting, and others. SAST scanners, also known as Source Code Analysis Tools, examine the source code without executing the application. They search for specific patterns in the code that are known to be vulnerable. Both types of scanners complement each other, as they can detect vulnerabilities that the other type might miss. Thereby it is crucial for stakeholders to incorporate both types of scanners into their infrastructure.
II-C Evolution of Large Language Models (LLM’s)
The evolution of Large Language Models (LLM’s) has significantly advanced their capabilities in vulnerability detection, with considerable potential for further growth. LLM’s have rapidly progressed, highlighted by key developments such as OpenAI’s GPT-4 [1] and Google Gemini 1.5 [24]. These models use token prediction to generate human-like text, demonstrating exceptional proficiency in various tasks related to language understanding. Recently, LLM has been applied to more interdisciplinary domains, such as code generation [22, 29, 12, 26], task planning and tool usage [19], LLM agent [10, 28], and even time-series prediction [15, 27, 14, 30],
II-D Integration of LLM’s in Cyber Security
Their ability to comprehend and generate text has made it possible to utilize LLM’s in software-related tasks. By recognizing and analyzing patterns in vast amounts of data, LLM’s have the potential to significantly enhance traditional vulnerability scanning techniques.
The use of LLM’s for both SAST (Static Application Security Testing) and DAST (Dynamic Application Security Testing) has proven feasible due to their advanced language processing capabilities. LLM’s can identify vulnerabilities that may be missed by traditional pattern matching and fuzzing techniques employed by conventional scanners.
II-E Privacy Concerns with LLM’s
Despite these advancements, the use of LLM’s raises significant privacy concerns, particularly regarding the handling and sharing of sensitive code data. Organizations must consider the potential risks associated with sending proprietary or confidential code to external LLM providers. These risks include data breaches due to vulnerabilities in network transmission, unauthorized access by LLM providers to sensitive intellectual property, compliance issues with data privacy regulations, and the loss of control over how the code is stored and processed once it leaves the organization’s infrastructure.
To mitigate these risks, organizations can use open-source LLM solutions that allow them to host these models locally. By hosting open-source LLM’s on their own infrastructure, organizations can reduce the risk of unauthorized access and data breaches while ensuring compliance with data privacy regulations and maintaining greater control over their sensitive data.
II-F Need for Up-to-Date Vulnerability Information
In addition to privacy concerns when using LLM’s for vulnerability detection, there are also issues regarding the absence of up-to-date vulnerability information. LLM’s are trained on extensive datasets compiled over a period leading up to the model’s creation. These datasets encompass a wide array of text and information from sources such as books, articles, and web content available up to that point. During training, these datasets must remain static to provide a stable sample of language patterns and context. Consequently, a cut-off date is introduced after which the training data is no longer updated, leading to the exclusion of recent events or information from the model’s knowledge.
The dynamic nature of the cyber security landscape requires constant awareness of the latest threats. New vulnerabilities, attacks, and coding errors are discovered frequently, and if identified after the cut-off date, they may not be included in the LLM’s knowledge base. While LLM’s primarily detect logic flaws rather than specific vulnerabilities, new types of attacks can introduce patterns or coding practices that may not be well represented in the training data of the model.
Therefore, relying solely on the static training dataset of the LLM may face limitations in detecting emerging security risks. This underscores the need for integrating up-to-date vulnerability information to enhance the effectiveness of LLM’s in vulnerability detection.
II-G Retrieval-Augmented Generation
To address this issue, we will introduce a new method of providing up-to-date vulnerability information to the LLM.
We will leverage Retrieval-Augmented Generation, a technique that enhances the capabilities of language models by integrating relevant external knowledge during the answer generation process. This external knowledge is stored as vector representations in databases, allowing it to be queried through similarity searches.
By conducting similarity searches, we can retrieve only the relevant external knowledge from these vector databases. This information is then used to enrich the LLM’s responses, ensuring that it incorporates relevant external knowledge into its generated answers.
III Related Paper
III-A Relevance of LLM’s in static vulnerability detection
Several papers have already demonstrated the relevance of utilizing LLM’s for static vulnerability detection, such as ”The Emergence of Large Language Models in Static Analysis” [21] and ”SkipAnalyzer” [16].
III-B Privacy concerns in LSAST
The paper ”Large Language Model for Vulnerability Detection” [31] emphasizes the importance of using locally hosted LLM’s to address privacy concerns associated with third-party LLM providers. However, the paper has not benchmarked these locally hostable open-source LLM’s.
III-C Importance of Finetuning in LSAST
Papers like ”SecureFalcon” [5] and ”Finetuning Large Language Models for Vulnerability Detection” [20] have shown that fine-tuning LLM’s can significantly enhance their vulnerability detection capabilities. These papers also recognize that the primary limitations of LLM’s in vulnerability detection extend beyond just model accuracy.
III-D Knowledge retrieval in LSAST
The paper ”LLM4Vuln” [23] highlights the critical role of robust knowledge retrieval when utilizing LLM’s for vulnerability scanning. It demonstrates LLM’s capabilities in scanning smart contract code and emphasizes that effective knowledge retrieval is crucial for building an LLM-based SAST scanner. However, the paper mainly explores various components of such a scanner, with the knowledge retrieval system playing a relatively minor role in their research.
Our paper proposes not only combining traditional SAST scanner results with LLM’s but also enhancing the knowledge retrieval aspect in multiple ways. We introduce new strategies for retrieving relevant vulnerability knowledge to support the LLM during the analysis phase.
Another paper, ”Vul-RAG” [4], implements a knowledge retrieval system using Retrieval-Augmented Generation (RAG). While similar to our approach in using code-based similarity search, it does not combine functionality-similarity search or assist conservative SAST scanners, opting instead for a standalone approach.
III-E Assisting conservative SAST scanners with LLM’s
Research on assisting conservative SAST scanners with LLM’s is sparse. ”The Hitchhiker’s Guide to Program Analysis” [11] recognizes the assistive role of LLM’s in static analysis and presents an initial approach for achieving synergy with LLM’s in static analysis. However, it focuses on a specific case study of UBI bugs (Use-Before- Initialization) in Linux kernels and does not leverage specific static analysis tools or support a wide range of programming languages as our approach does. In our work we evaluate LSAST on a wide range of programming languages including javascript, php, java and others.
Another relevant paper, ”LLM-Assisted Static Analysis for Detecting Security Vulnerabilities” [13] leverages LLM’s to complement traditional static analysis techniques, focusing on using CodeQL for detailed code querying and analysis. This approach does not directly integrate SAST scanner outputs into the LLM prompt or implement knowledge retrieval systems as described in LSAST, but it has the potential to be integrated with our research in the future.
@@ IV Methodology
IV-A Combining conservative SAST scanners with LLM’s
The main aspect of LSAST is the combination of a traditional SAST scanner with a large language model (LLM). To implement LSAST, it is crucial to choose an appropriate SAST scanning tool. We selected the open-source SAST scanner Bearer due to its fast response times and its ability to output data in various formats, which is useful for integrating the results into the prompt for the LLM.
@@ IV-A Combining conservative SAST scanners with LLM’s
In Listing 1, you can see an example of a simplified output from a single Bearer scan finding. This output includes crucial details, such as the type of vulnerability and the affected lines.
In Listing 2, you can see how we formatted the output of a Bearer scan.
@@
These results can now be integrated into the prompt for the LLM. Additionally, we need to provide the LLM with the target code we want to analyze. After importing the code file as a string, we adjust each line of the code to include line number information.
In Listing 3, you can observe the appearance of the target code string after the necessary adjustments. Including line number information in the code ensures that the SAST response, which also contains line number information, can be interpreted effectively by the LLM. Failure to include line numbers could result in the loss of the LLM’s ability to utilize the SAST scanner results comprehensively, potentially leading to inaccuracies in the LLM’s own findings regarding line-specific vulnerabilities.
@@
Listing 4 presents the complete prompt, encompassing both the target code and the results from the Bearer scan. This integrated prompt is now ready for direct querying by the LLM.
@@ IV-B Gathering Vulnerability Reports
To ensure our queries include the most current vulnerability information, we’ve implemented a Knowledge Retrieval System. This system aggregates vulnerability reports from the past year, offering insights into newly exploitable vulnerabilities that may not be covered in the LLM’s training data.
To begin, we retrieve these reports and prepare them for a similarity search. We utilize the Hacktivity API from [7] for gathering vulnerability data. HackerOne, a leading Bug Bounty Platform, hosts programs where companies incentivize security researchers to uncover vulnerabilities in their systems. Some findings from these programs are openly disclosed on HackerOne and accessible via their API. A collection of vulnerability reports is invaluable for LLM-supported DAST or SAST scanners, particularly in scenarios involving both Black-Box (application-only access) and White-Box (full code access) approaches.
Our efforts have yielded 873 vulnerability reports in JSON format, all dated within the past year and containing software code details.
In Listing 5, you can find a simplified example of a single vulnerability report. This report includes the title, severity rating (ranging from low to critical), CVE IDs, CWE ID and title, and most importantly, vulnerability information. The vulnerability information section provides a detailed description of the vulnerability along with relevant code examples.
@@
IV-C Preparing Vulnerability Reports for Knowledge Retrieval Process
To deliver these vulnerability reports effectively to the LLM, we employ a Retrieval-augmented System, which enhances the LLM’s internal representation of information. Given that supplying all 873 vulnerability reports could potentially overwhelm and distract the LLM, our approach focuses on providing only the most relevant reports for the code that has to be scanned.
Various methods can determine which reports are most relevant for the target code.
LLM4Vuln [23] introduced a method that involves a similarity search based on the functionality of the code.

### v2

@@ Abstract
The current cybersecurity landscape is increasingly complex, with traditional Static Application Security Testing (SAST) tools struggling to capture complex and emerging vulnerabilities due to their reliance on rule-based matching. Meanwhile, Large Language Models (LLMs) have demonstrated powerful code analysis capabilities, but their static training data and privacy risks limit their effectiveness. To overcome the limitations of both approaches, we propose LSAST, a novel approach that integrates LLMs with SAST scanners to enhance vulnerability detection. LSAST leverages a locally hostable LLM, combined with a state-of-the-art knowledge retrieval system, to provide up-to-date vulnerability insights without compromising data privacy. We set a new benchmark for static vulnerability analysis, offering a robust, privacy-conscious solution that bridges the gap between traditional scanners and advanced AI-driven analysis. Our evaluation demonstrates that incorporating SAST results into LLM analysis significantly improves detection accuracy, identifying vulnerabilities missed by conventional methods.
@@ I Introduction
With the global loss of cybercrime projected to increase from \$8.15 trillion in 2023 to \$13.82 trillion in 2028 [1] and hackers launching an average of 26,000 attacks daily—one every three seconds [2]—cybersecurity is more critical than ever in the digital age. As the IT landscape grows, cyber threats have evolved from simple attacks to sophisticated and persistent threats, necessitating the implementation of comprehensive cybersecurity measures.
One widely adopted measure is security scanning[3, 4] to find out the potential vulnerabilities [5, 6], which plays a crucial role in identifying potential risks within applications and networks. However, cybersecurity threats are continuously evolving, and new vulnerabilities emerge rapidly, making it increasingly difficult for traditional security detection tools to keep pace. Static Application Security Testing (SAST)[7], a conventional static analysis tool, is commonly used for code security testing, but its rule-based pattern matching approach limits its effectiveness in detecting complex vulnerabilities. Large Language Models (LLMs) [8, 9] have demonstrated significant potential in code analysis, particularly with their abilities to understand complex contexts and detect diverse vulnerabilities. By retrieving the latest vulnerability reports, LLMs can learn new types of vulnerabilities in real-time without being constrained by static training data[10, 11, 12]. Unlike traditional pattern matching and fuzzing techniques, LLMs can identify vulnerabilities that may be overlooked by conventional scanners[13, 14, 15].
However, since LLMs’ training data is not updated after the initial training, they struggle to promptly recognize newly emerging security threats. This limitation poses substantial security risks when using LLMs for vulnerability scanning, as effective scanning depends on accessing to up-to-date information on the latest vulnerabilities to adequately address emerging threats.
@@
Additionally, using LLMs for vulnerability scanning introduces further challenges, particularly when code is transmitted to third-party LLM providers via API endpoints. This practice can expose sensitive code, leading to potential privacy and security risks, including unauthorized access by malicious actors. Moreover, there is a risk of the code being repurposed for future model training, making it vulnerable to training data extraction attacks. Addressing these concerns is essential to prevent the exploitation of potential vulnerabilities in the code.
In this paper, we aim to tackle these challenges head-on. We
  • enhance LLMs’ capabilities in vulnerability scanning by integrating static code scanning results with LLMs, allowing them to detect vulnerabilities that would be missed by traditional methods alone. This demonstrates the powerful synergy achieved by combining these technologies.
  • address the limitation of LLMs in accessing up-to-date vulnerability information by introducing novel approaches for knowledge retrieval systems. Our combined and abstraction-based approaches outperform existing methods and set new standards for building and enhancing knowledge retrieval systems for static vulnerability scanning.
  • demonstrate that knowledge retrieval systems can be counterproductive if not carefully designed, as irrelevant reports can distract the LLM’s focus.
  • mitigate privacy and security risks by utilizing open-source LLMs, showing that LLMs can be leveraged for vulnerability scanning without compromising these critical areas.
  • validate in real-world scenarios that LSAST scanners can identify vulnerabilities overlooked by traditional scanners, establishing their importance in comprehensive security scanning processes.
The rest of the paper is organized as follows. Section II presents the background of security scanners and vulnerability information. Section III reviews the related work on security testing tools, LLMs in vulnerability detection and Knowledge retrieval. Section IV gives the LSATA and details the optimization based on SAST. Section V presents our experimental results and findings. Finally, Section VI concludes the paper.
@@ II Background
In this Section, we will first introduce different kinds of existing security scanners, followed by Retrieval-Augmented Generation (RAG).
II-A Different kinds of security scanners
@@
Predominantly, security scanning can be classified into two categories: DAST (Dynamic Application Security Testing) and SAST (Static Application Security Testing)[16]. DAST scanners treat the system or application as a black-box, meaning they operate without knowledge of the internal workings. They identify weaknesses by interacting with the system or application during runtime. By sending specially crafted requests and analyzing the responses, DAST scanners mimic threat actors attempting various types of attacks, such as SQL Injection, Cross-Site Scripting, and others. SAST scanners, also known as Source Code Analysis Tools, examine the source code without executing the application. They search for specific patterns in the code that are known to be vulnerable. Both types of scanners complement each other, as they can detect vulnerabilities that the other type might miss. Thereby it is crucial for stakeholders to incorporate both types of scanners into their infrastructure.
II-B Retrieval-Augmented Generation
As the cybersecurity landscape evolves rapidly, staying updated with the latest vulnerabilities and attack vectors is crucial for effective defense. New security threats, vulnerabilities, and coding errors are discovered frequently, and traditional approaches to static analysis struggle to keep pace with these emerging risks. LLMs, while powerful in detecting logic flaws and complex vulnerabilities, are limited by their static training data, which often does not include the most recent security developments.
To address this limitation, we leverage Retrieval-Augmented Generation (RAG), a technique that dynamically enhances LLM capabilities by integrating relevant external knowledge during the response generation process. RAG bridges the gap between static training data and the need for real-time awareness of the latest vulnerabilities. In RAG, external knowledge is stored as vector representations within specialized databases. During the LLM’s processing, similarity searches are conducted within these databases to retrieve the most relevant information. This approach allows the LLM to draw upon a continuously updated knowledge base, incorporating fresh insights and vulnerability reports that are not part of the original training data.
III Related Work
This section reviews existing research on vulnerability detection methods, highlighting the limitations of traditional approaches, the emerging role of LLMs, and the integration of knowledge retrieval systems.
III-A Limitations of Traditional Methods
Traditional static and dynamic application security testing tools, such as SAST and DAST, have long been the cornerstone of vulnerability detection. However, several studies have identified significant limitations in their effectiveness, particularly when dealing with complex and context-specific vulnerabilities. For example, SAST tools often rely on predefined patterns and rules, which can result in high false-positive rates and an inability to detect novel or context-dependent vulnerabilities. This issue is underscored in [17] and [18]. Attempts to integrate machine learning or heuristic-based approaches into SAST and DAST have been explored, yet they often lack the depth of understanding needed for sophisticated vulnerability detection. Li et al. [19] recognize the potential of combining LLMs with traditional static analysis but focus narrowly on specific bugs, such as Use-Before-Initialization (UBI) in Linux kernels, and do not fully leverage specific static analysis tools or support a wide range of programming languages. These gaps have led researchers to explore the use of LLMs as a way to enhance the effectiveness of traditional static analysis tools.
III-B Other Aspect of Vulnerability
Prior work has explored various aspects of security and privacy vulnerabilities in machine learning models. The study in [6] addresses model inversion attacks (MIAI), proposing a novel black-box approach to infer sensitive information about the training data while highlighting the disparate vulnerability impact on specific demographic groups. However, it focuses solely on privacy risks in machine learning models and does not address static code analysis or vulnerability detection in software applications. Meanwhile, [5] investigates the security implications of deploying efficient, on-device deep learning models, showing that these models can increase adversarial vulnerabilities in their full-scale counterparts. Although it introduces a defense mechanism to mitigate these risks, the work is limited to adversarial attacks and does not extend to static vulnerability detection. In contrast, our proposed LSAST framework addresses these limitations by integrating Large Language Models (LLMs) with traditional Static Application Security Testing (SAST) tools, enhancing vulnerability detection in software applications while maintaining data privacy and providing up-to-date insights, which are not covered in these previous works.
III-C LLMs in Vulnerability Detection
LLMs have emerged as a powerful tool in the field of vulnerability detection, offering a significant upgrade over traditional scanners due to their ability to understand code in a more context-aware manner. Studies such as [20] and [21] have demonstrated that LLMs can effectively identify vulnerabilities that static tools often miss, thanks to their deep learning-based understanding of code structure and semantics. Despite these advantages, LLMs face their own set of challenges. One major concern is privacy, particularly when using third-party LLM services that require codes to be sent externally for analysis. Zhou et al. [22] emphasize the importance of using locally hosted LLMs to address these privacy concerns, although this approach has not yet been fully benchmarked for its effectiveness in static vulnerability scanning. Another critical challenge is the static nature of LLM training data. While fine-tuning techniques highlighted in [23] can improve model accuracy, LLMs still struggle to adapt to the latest security threats without frequent updates. These limitations underscore the need for integrating real-time knowledge retrieval systems to dynamically supply LLMs with up-to-date information, enhancing their detection capabilities.
III-D Enhancing LLMs with Knowledge Retrieval and SAST Integration
Recent research has explored the use of knowledge retrieval systems to address the limitations of LLMs, particularly their reliance on static training data. LLM4Vuln[24] highlights the critical role of robust knowledge retrieval in enhancing LLM-based vulnerability scanning, demonstrating that such systems can provide LLMs with current data and improve their scanning capabilities. However, this study primarily focuses on scanning smart contracts and does not deeply explore the integration of SAST results with knowledge retrieval. Du et al. [25] implement a RAG system to enhance LLMs with real-time data, using code-based similarity searches. While this approach shares similarities with our method, it does not combine functionality-similarity search or directly support SAST scanners, opting instead for a standalone approach without leveraging SAST outputs. Efforts to directly assist SAST with LLMs have been limited, with research often focusing on specific case studies rather than broader implementations. Li et al. [18] use CodeQL for detailed code querying and analysis but does not integrate SAST scanner outputs into the LLM prompt or employ advanced knowledge retrieval systems as we propose.
@@ IV Methodology
In this section, we will first introduce the idea of combining conservative SAST scanner with LLMs in Section IV-A, and then gather vulnerability reports in Section IV-B and prepare them ready for the knowledge retrieval process in Section IV-C. Then, we design and test seven different knowledge retrieval methods in Section IV-D.
IV-A Combining conservative SAST scanners with LLMs
To implement LSAST, it is crucial to choose an appropriate SAST scanning tool. We selected the open-source SAST scanner Bearer due to its fast response times and its ability to output data in various formats, which is useful for integrating the results into the prompt for the LLM.
An example of a simplified output from a single Bearer scan finding is shown in Listing 1, with crucial details, such as the type of vulnerability and the affected lines. To better integrate these findings into the prompt for the LLM, we format the output in the way shown in Listing 2. Additionally, we need to provide the LLM with the target code we want to analyze. After importing the code file as a string, we adjust each line of the code to include line number information.
@@ IV-A Combining conservative SAST scanners with LLMs
Target code string after adjustments is demonstrated in Listing 3. Including line number information in the code ensures that the SAST response, which also includes line numbers, can be accurately interpreted by the LLM. Omitting line numbers may hinder the LLM’s ability to fully utilize SAST scanner results, potentially leading to inaccuracies in the LLMs own findings regarding line-specific vulnerabilities.
Furthermore, Listing 4 presents the complete prompt, encompassing both the target code and the results from the Bearer scan. This integrated prompt, which includes detailed instructions, is now ready for direct querying by the LLM.
@@ IV-B Gathering Vulnerability Reports
To ensure our queries include the most current vulnerability information, we’ve implemented a Knowledge Retrieval System. This system aggregates vulnerability reports from the past year, offering insights into newly exploitable vulnerabilities that may not be covered in the LLMs training data.
@@
To begin, we retrieve these reports and prepare them for a similarity search. We utilize the Hacktivity API [^1] HackerOne, (https://hackerone.com/hacktivity/), a leading Bug Bounty Platform, hosts programs where companies incentivize security researchers to uncover vulnerabilities in their systems. Some findings from these programs are openly disclosed on HackerOne and accessible via their API. A collection of vulnerability reports is invaluable for LLM-supported DAST or SAST scanners, particularly in scenarios involving both Black-Box (application-only access) and White-Box (full code access) approaches. for gathering vulnerability data. Our efforts have yielded 873 vulnerability reports in JSON format, all dated within the past year and containing software code details.
We provide a simplified example of a single vulnerability report in Listing 5. This report includes the title, severity rating (ranging from low to critical), CVE IDs, CWE IDs and title, and vulnerability information. The vulnerability information section provides a detailed description of the vulnerability along with relevant code examples.
IV-C Preparing Vulnerability Reports for Knowledge Retrieval
To effectively deliver vulnerability reports to the LLM, we utilize a retrieval-augmented system that combines SAST scan results with the target code. This integration enhances the LLMs understanding by providing a more complete context for analysis. Instead of overwhelming the LLM with all 873 available vulnerability reports, which could lead to distractions and reduced performance, our approach selectively retrieves only the most relevant reports that match the specific characteristics of the code being scanned. This targeted delivery ensures that the LLM focuses on pertinent information, thereby improving its accuracy in identifying new vulnerabilities.